# Learning Informative Latent Representation for Quantum State Tomography

### v1

@@ 3.3 Learning an informative latent representation for QST with imperfect measurement data
A density matrix $\rho$ can be generated from a lower triangular matrix $\rho_{L}\in\mathbb{C}^{d\times d}$ according to $\rho=\rho_{L}\rho_{L}^{\dagger}/{\textup{Tr}(\rho_{L}\rho_{L}^{\dagger})}$, satisfying (i) $\rho=\rho^{\dagger}$, (ii) $\textup{Tr}(\rho)=1$, and (iii) $\rho\geq 0$. Conversely, any $\rho$ can be decomposed using the Cholesky decomposition [16], i.e., $\rho_{L}\rho_{L}^{\dagger}=\rho$. Furthermore, $\rho_{L}$ can be transformed into a real vector $\bm{\nu}\in\mathbb{R}^{d^{2}\times 1}$ by concatenating its real and imaginary parts. Hence, retrieving a real vector $\bm{\nu}$ is equivalent to recovering the density matrix $\rho$.
4 Experiments
4.1 Implementation details
Quantum states and measurements. We consider both pure and mixed quantum states, where pure states can be generated by performing random unitary matrices on a pure state following the Haar metric [7] and mixed states can be generated from the Ginibre ensembles [13] using the Hilbert-Schmidt metric [33]. We adopt the cube measurement [8], which contains $M=6^{n}$ operators for $n$-qubit systems, which is over-complete compared to the minimum number (i.e., $4^{n}$) of linearly independent measurements for informationally complete QST. Please refer to Appendix A for detailed information. Measured data collection. 100,000 quantum states are randomly generated for both pure and mixed states, with 95,000 samples used for training and 5,000 for testing. Provided with states and measurement operators, we obtain true probabilities according to Eq. (1) and simulate the observation process on desktop CPUs using multinomial distribution functions with true probabilities and total copies. Upon collecting the concurrence, the frequencies are obtained according to Eq. (2). Neural network architecture. The ILR-QST model is implemented on PyTorch with the following specifications: 1) Encoder: The transformer encoder consists of 8 layers with 16 attention heads (each with a dimension of 32) and 256 hidden units. The input to the encoder is a sequence of the feature latent $\tilde{E}_{i}$. The encoder outputs a sequence of latent features with the same dimension as the input. 2) Decoder: The parameters of the transformer model in the frequency encoder and state encoder are the same as those in the encoder, except for the number of layers. The frequency decoder employs one layer of the transformer block, whereas the state decoder consists of four transformer layers for reconstructing $\bm{\nu}$ (or equivalently $\rho$) and one transformer layer for predicting $\bm{\mu}$.
Training settings. The total copies are allocated into several groups, with each containing $d$ measurement operators. We define the number of copies involved in $d$ operators as $N_{t}$ to denote an average value. Hereafter, we use $m$ and $N_{t}$ to represent two imperfect factors in practical QST applications. Given a fixed $N_{t}$ (e.g., $N_{t}=100$), the pre-train process is performed on a large number of imperfect samples, with $m$ randomly sampled from its predefined set, e.g., $\{0,4,8,...\}$, while the QST process is then fine-tuned on a specific $m$ separately. In the training process, we utilize the default Adam optimizer, a batch size of 256, and an initial learning rate of $35E-$ 3. To improve convergence and reduce the risk of overfitting, we employ a strategy of cosine learning rate decay. The model is trained for 500 epochs with a warm-up strategy for the first 20 epochs, during which the learning rate is gradually increased from 0 to $35E-$ 3.
4.2 Reconstructing density matrices
(a) 2-qubit pure states
(b) 4-qubit pure states
Figure 4: Comparison results of reconstructing density matrices for 2-qubit states ($N_{t}=100$).
Here, we implement the FCN [29], LRE [35], and MLE [21] for comparison. To conduct an overall comparison, we compare the number of parameters, and inference time (GPU/CPU) for the proposed ILR and other methods. The simulation platform consists of an 8-core Intel(R) Xeon(R) W-2145 CPU @ 3.70GHz, 64G DDR4 memory size, and Quadro RTX 4000 with 8G. For inference time (CPU/GPU), we conduct 5,000 times of reconstructing density matrix and report the average time per inference in Table 1.
Table 1: Comparison of parameters and inference time for ILR and other methods used in Figure 4. Infer-CPU/GPU denotes the inference process performed on CPU/GPU. Despite the need for training and parameters in NN approaches, they consistently demonstrate superior performance to traditional methods. Notably, our ILR model outperformed FCN approaches in terms of performance with comparable parameters and inference time.
2 qubit 4 qubit
Methods Infidelity Params Infer-CPU Infer-GPU Infidelity Params Infer-CPU Infer-GPU
LRE 3.54E-2 N/A 4.57ms N/A 9.93E-2 N/A 671ms N/A
MLE 3.56E-2 N/A 86.6ms N/A 8.14E-2 N/A 793ms N/A
FCN 1.80E-3 211K 5.42ms 1.30ms 6.25E-2 595K 20.35ms 1.88ms
ILR 1.85E-4 218K 6.92ms 1.38ms 1.28E-2 257k 23.71ms 2.23ms
The numerical results for 2-qubit and 4-qubit pure states are summarized in Figure 4, where the log of the infidelity of the four methods increases with the number of masked operators, which is in agreement with theoretical expectations. The proposed approach (ILR) outperforms FCN and exhibits a superiority over LRE and MLE. Investigation of different number of copies ($N_{t}=10$ and $N_{t}=1000$) and 2-qubit mixed states are summarized in Appendix C. Those results suggest that the ILR is effective in reconstructing density matrices from imperfect measurement data.
We also perform experiments on IBM quantum machine ibmq_manila and results for 50 samples are provided in Table 2. Note that our ILR model is trained using simulated data that do not consider the noise model in real quantum devices. The results show the excellent robustness of our ILR model in dealing with experimental data, compared with LRE and the IBM-built method.
Table 2: Statistical fidelities (from 50 samples) of ILR (model trained in Figure 4) and other methods on IBM quantum computer. Here a higher value of fidelity means a better performance.
Simulation data Experimental data (ibmq_manila)
ILR LRE ILR LRE IBM-built in
Min 0.999463 0.911296 0.930446 0.833402 0.819149
Max 0.999982 0.995680 0.993310 0.960062 0.957353
Mean 0.999828 0.972375 0.973660 0.901905 0.892300
Variance 2.42E-08 6.01E-04 3.16E-04 1.06E-03 1.15E-03
4.3 Predicting quantum properties
Refer to captionRefer to captionRefer to caption
(a) 2-qubit pure states
Refer to captionRefer to captionRefer to caption
(b) 4-qubit pure states
Figure 5: Results of predicting properties for pure states ($N_{t}=100$). In ILR-B, FCN, MLE, and LRE (dash lines), quantum properties are calculated from the reconstructed density matrices.
Apart from directly using ILR to predict quantum properties, we introduce a baseline with properties calculated from density matrices using ILR in Subsection 4.2 (termed ILR-B). Please refer to Appendix E for their detailed comparsion. Then, we also summarize the results of calculating properties from reconstructed density matrices using FCN, LRE, and MLE. To distinguish them from ILR, dash lines are utilized for the above four methods. The results for 2-qubit and 4-qubit states at $N_{t}=100$ are provided in Figure 5, with results of 2-qubit mixed states presented in Appendix D. The proposed ILR directly predicts properties with the highest accuracy, with great superiority over the indirect methods that first recover the density matrix using traditional algorithms and then calculate properties. Those results demonstrate the informative representation has great potential to deduce a partial description of quantum states under imperfect conditions.
To demonstrate the applicability of ILR to high-dimensional states, we consider physical states with specific patterns, i.e., locally rotated GHZ states of the form $\otimes_{i=1}^{n}U_{i}|\psi_{GHZ}\rangle$ with $|\psi_{GHZ}\rangle=\frac{1}{\sqrt{2}}(|00...00\rangle+|11...11\rangle)$, where $U_{i}$ denotes a random unitary transformation in the $i$-th qubit. We also consider locally rotated W states of the form $\otimes_{i=1}^{n}U_{i}|\psi_{W}\rangle$ with $|\psi_{W}\rangle=\frac{1}{\sqrt{n}}(|100...00\rangle+|010...00\rangle+...+|000...10\rangle+|000...01\rangle)$. We utilize two-qubit Pauli measurements on nearest-neighbor qubits, with $(n-1)*9*4$ types of measurement [45]. For 8 qubits, we utilize 9000 states for training and 1000 states for testing. For 12 qubits, we utilize 900 states for training and 100 states for testing. From Table 3, our method has good stability in predicting the properties of quantum states. This demonstrates that ILR has great potential to deduce a partial description of quantum states under imperfect conditions.
Table 3: MSE of predicting properties on 8/12-qubit GHZ and W states with $N_{t}=100$. Mask Level (1,2,3,4): (0, 16, 24, 32) out of 63 for 8-qubit; (0, 24, 36, 48) out of 99 for 12-qubit.
Mask Level 8-qubit 12-qubit 8-qubit 12-qubit
5 Conclusion
In this paper, we investigated QST with imperfect measurement data, i.e., a few copies of identical states to approximate true probabilities or incomplete measurements that theoretically fail to specify a unique solution. Inspired by the architecture of an autoencoder that can learn latent representation from imperfect data, we designed a transformer-based autoencoder to abstract an informative latent representation from imperfect measurement data. Drawing the similarity of incompleteness and mask, we designed a masked autoencoder to map the raw measured frequencies into high-quality probabilities. By pre-training the masked autoencoder with a large number of samples obtained from ill-posed scenarios, a highly informative latent representation has been extracted and utilized to efficiently reconstruct quantum states from imperfect measured outcomes. The numerical simulations and experimental implementations demonstrate that ILR provides a robust and efficient solution for QST with imperfect measurement data. Our future work will focus on neural network-based quantum process tomography.
Appendix A Quantum states, measurements and properties
Quantum states. In this work, we consider both pure states and mixed states. To generate pure random states, we utilize the Haar measure [30] to generate random unitary matrices. Let $\mathbb{U}^{d}$ be the set of all $d$-dimensional unitary operators, i.e., any $U\in\mathbb{U}^{d}$ satisfies $UU^{\dagger}=U^{\dagger}U=\mathbb{I}_{d}$. Owing to its invariance under group multiplication (i.e., any region of $\mathcal{U}^{d}$ carries the same weight in a group average) [30], the Haar metric is utilized as a probability measure on a compact group. This property enables us to generate random unitary transformations that are further utilized to generate random pure states as
$\rho_{haar}=U|\psi_{0}\rangle\langle\psi_{0}|U^{\dagger},$ (3)
where $|\psi_{0}\rangle$ is a fixed pure state. For mixed states, we consider the random matrix from the Ginibre ensembles  [13] given by $\rho_{G}=\mathcal{N}(0,1)_{d}+\textup{i}\mathcal{N}(0,1)_{d}$, where $\mathcal{N}(0,1)_{d}$ represents random normal distributions of size $d\times d$ with zero mean and unity variance. Random density matrices using the Hilbert-Schmidt metric are given by [33]
$\rho_{ginibre}=\frac{\rho_{G}\rho_{G}^{\dagger}}{\textup{Tr}(\rho_{G}\rho_{G}^{\dagger})}.$ (4)
Quantum measurements. As for measurements, we consider tensor products of Pauli matrices, which is also called cube measurement [8]. Denote the Pauli matrices as $\sigma=(\sigma_{x},\sigma_{y},\sigma_{z})$, with
$\sigma_{x}=\left[\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right],\sigma_{y}=\left[\begin{array}[]{cc}0&-\rm{i}\\ \rm{i}&0\end{array}\right],\sigma_{z}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right].$ (5)
Let $|H\rangle$ and $|V\rangle$ be eigen-vectors of $\sigma_{z}$, $|L\rangle$ and $|R\rangle$ be eigen-vectors of $\sigma_{y}$, and $|D\rangle$ and $|A\rangle$ be eigen-vectors of $\sigma_{x}$. The POVM elements for one-qubit Pauli measurement are finally given as $P_{pauli}=\{|H\rangle\langle H|,|V\rangle\langle V|,|L\rangle\langle L|,|R\rangle\langle R|,|A\rangle\langle A|,|D\rangle\langle D|\}$. Generalizing the 1-qubit Pauli measurement to $n$-qubit using tensor product, we obtain the following POVMs
$P=P_{1}\otimes P_{2}\cdots\otimes P_{n},\quad\text{with}\quad P_{i}\in P_{pauli}.$ (6)
For $n$-qubits, there are $6^{n}$ POVMs involved in the cube measurement. Typically, the $6^{n}$ measurement operators can be arranged into several groups, with each group containing $d=2^{n}$ elements [1]. For example, $|HH\rangle\langle HH|$,$|HV\rangle\langle HV|$,$|VH\rangle\langle VH|$, and $|VV\rangle\langle VV|$ can be allocated into one group. These four POVM elements actually compose a 2-qubit detector that can be experimentally realized on quantum devices.
Table 4: Descriptions for quantum properties.
Name Definition Range
Purity $\textup{Tr}(\rho^{2})$ $[0,1]$
Entropy $\textbf{S}(\rho):=-\textup{Tr}(\rho\ln(\rho))$ $[0,\ln(d)]$
Coherence $S(\rho_{diag})-S(\rho)$ $[0,\ln(d)]$
Entanglement
entropy $S(\textup{Tr}_{A}(\rho))$ $[0,\ln(d)/2]$
Negativity $\frac{||\rho^{\Gamma_{A}}||_{1}-1}{2}$ $[0,0.5]$
Quantum properties. In this work, we consider six properties that are widely used in quantum information processing tasks. Pure states have a fixed purity of 1 and a fixed entropy of 0, which can be proven by simple calculation from the equations in Tabel 4. Coherence is commonly measured by the sum of squared values of non-diagonal elements among the density matrix, i.e., $\sum_{i,j,i\neq j}|\rho_{ij}|$, which can be extremely large for high-dimensional cases. Instead, we consider relative coherence regarding the entropy [2]. The von Neumann entanglement entropy is defined as the von Neumann entropy of either of its reduced states since they are of the same value, which can be proved from Schmidt’s decomposition of the state with respect to the bipartition. Concurrence is originally defined for 2-qubit mixed states, with $\{\lambda_{i}\}$ being the eigenvalues (decreasing) of the Hermitian matrix $R=\sqrt{\sqrt{\rho}\rho_{y}\sqrt{\rho}}$ with $\rho_{y}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$, with the complex conjugation $*{}^{*}$ taken in the eigenbasis of the Pauli matrix $\sigma_{z}$. A generalized version of concurrence for multiparticle pure states in arbitrary dimension is defined as $\sqrt{2(1-\textup{Tr}(\textup{Tr}_{A}(\rho)^{2}))}$. Among the six properties in Tabel 4, the latter three properties are designed to evaluate the entanglement of quantum states. Specifically, their values achieve the lowest value of zero for separable states and achieve the highest value for maximally entangled states, e.g., Greenberger–Horne–Zeilinger states.
Appendix B Additional information about training strategies
Algorithm descriptions. The algorithm descriptions for the pre-training process and the QST process are presented in Algorithm 1 and Algorithm 2, respectively.

### v2

@@ III-C Learning an informative latent representation for QST with imperfect measurement data
A density matrix $\rho$ can be generated from a lower triangular matrix $\rho_{L}\in\mathbb{C}^{d\times d}$ according to $\rho=\rho_{L}\rho_{L}^{\dagger}/{\textup{Tr}(\rho_{L}\rho_{L}^{\dagger})}$. The matrices satisfy three requirements: (i) $\rho=\rho^{\dagger}$, (ii) $\textup{Tr}(\rho)=1$, and (iii) $\rho\geq 0$. Conversely, any $\rho$ can be decomposed using the Cholesky decomposition [47], i.e., $\rho_{L}\rho_{L}^{\dagger}=\rho$. Furthermore, $\rho_{L}$ can be transformed into a real vector $\bm{\nu}\in\mathbb{R}^{d^{2}\times 1}$ by concatenating its real and imaginary parts. Hence, retrieving a real vector $\bm{\nu}$ is equivalent to recovering the density matrix $\rho$.
III-D Algorithm description
The pre-training process is first implemented to obtain a highly representative representation of quantum states. Then, complete training of the QST process is achieved by keeping the parameters of the transformer-encoder fixed and fine-tuning the parameters of the state decoder. This structure enables the preservation of the information that is learned from the pre-trained model, and the latent representation enables the state decoder to reconstruct quantum states with improved performance. To simulate different incomplete measurement scenarios, we introduce two ways for pre-training: i) Separate, fixed mask values are utilized for all samples, and ii) Unified, where mask values are randomly selected for each training batch. For both cases, the same number of samples are used for training, while the learned representation regarding MSE can be different for training, which can be demonstrated in Fig. 4 of Section IV-B. Similarly, the QST process has the same two training options. The algorithm descriptions for the pre-training process and the QST process are presented in Algorithm 1 and Algorithm 2 (in Appendix G), respectively. The difference in this latent representation also influences the individual QST tasks; see detail in Fig. 4.
Now, there are four different strategies for training the whole model, including S2S (separate pre-training with separate tomography), S2U (separate pre-training with unified tomography), U2S (unified pre-training with separate tomography), and U2U (unified pre-training with unified tomography). We conduct a comprehensive comparison of 2-qubit pure states and determine U2S as a good strategy that balances accuracy and efficiency.